Operations Report – April 20th

Crew 297 Operations Report 20-04-2024

SOL: 6

Name of person filing report: Matt Lynch (Engineer)

Non-nominal systems: Suit 2 (stuck valve), Suit 9 (broken visor), Suit 11 (poor battery life)

Notes on non-nominal systems: The issues with Suit 2 and 9 were already noted before our mission. Suit 2 & 11 repairs are ongoing. The issue with Radio two has been fixed with a new headset.


Opportunity rover used:

Hours: 181.8

Beginning charge: 100%

Ending charge: 48%

Currently charging: No

Curiosity rover used:

Hours: 278.7

Beginning charge: 100%

Ending charge: 39%

Currently charging: No

General notes on rovers: Both rovers seats have been partly repaired with reflective silver metallic tape.

Summary of Hab operations:

WATER USE: 25.4 gallons

Water (static tank): 284.2 gallons

Static tank pipe heater (on or off): off

Static tank heater (On or off): off

Toilet tank emptied: no

Summary of internet: Looking up recipes, downloading entertainment, report formatting methods, catching up on Earthly news

Summary of suits and radios: The battery on suit 6 seemed to drain faster than the others today. While still currently within acceptable performance this should be monitored.

Radios 2,3,4,5, and 6 were used.

EVA Suits 1,4,5,and 6 were used.

Summary of GreenHab operations:

WATER USE: 4 gallons

Heater: Yes

Supplemental light: N/A

Harvest: 13g of cherry tomato, 66g of red cabbage, 147g of radishes, and 177g of kale, 23g of green onion

Summary of ScienceDome operations: NA

Summary of RAM operations: (Tools used, work done): Tools from the RAM have been continually used in suit repair.

Summary of any observatory issues: NA

Copyright © The Mars Society. All rights reserved. | Main Site